# Adversaries With Incentives: A Strategic Alternative to Adversarial Robustness

### v1

@@ 4.1 Adversarial attacks as worst-case strategic responses
In Eq. (3), the form for $\delta^{\mathrm{adv}}$ implies that the adversary benefits from any erroneous prediction; thus, an adversarial opponent is in fact a strategic opponent with utility $u=u^{\mathrm{adv}}\triangleq\ell$ plugged into Eq. (4). In general, however, strategic opponents can have other, non-adversarial utilities. Assuming utilities are normalized, let $\Lambda=[0,1]^{K\times K}$ denote the universal set of all utility functions. Then for an adversary, since $u^{\mathrm{adv}}_{y{\hat{y}}}=1$ if $y\neq{\hat{y}}$ and 0 otherwise, our first observation is that Eq. (3) can be rewritten as:
@@
Figure 3: (Left:) Accuracy of an adversarially trained model on varying opponents: adversarial, clean, and strategic (1-hots and 3-hots), vs. clean accuracy of a clean model. Strategic attacks are effective, but less severe than adversarial—a gap which strategic training aims to capitalize on (right).
@@ 4.2 The price of adversarial robustness against non-adversarial opponents
Fig. 3 shows the performance of an adversarially trained classifier ($f_{\mathrm{adv}}$) evaluated against different opponents: adversarial ($\mathrm{adv}$), strategic ($\mathrm{1{\text{-}}hot}$ and $\mathrm{3{\text{-}}hot}$), and no opponent ($\mathrm{clean}$). Against $\mathrm{adv}$ (dark red), for which $f_{\mathrm{adv}}$ is consistent in training, accuracy is $45.5$. This increases to $79.8$ (purple) in $\mathrm{clean}$. This is an improvement—but in comparison, a model trained on clean data ($f_{\mathrm{cln}}$) achieves $93.1$ clean accuracy. The price $f_{\mathrm{adv}}$ pays for being conservative is therefore quite steep ($-13.3$). Under our utilitarian perspective, this large gap is the result of assuming that all attacks are valuable for the opponent, when in effect none are.
@@
Next, consider how the same adversarially trained model fairs against strategic opponents. Fig. 3 also shows the distribution of strategic accuracies of $f_{\mathrm{adv}}$ against all $k=1$ (dark blue) and a large sample of $k=3$ (light blue) row-wise $k$-hot opponents. Results show that strategic attacks are indeed generally weaker than adversarial attacks, with average values of $69.9\mathrm{1{\text{-}}hot}(f_{\mathrm{adv}})=$ 69.9 and $58.5\mathrm{3{\text{-}}hot}(f_{\mathrm{adv}})=$ 58.5. These large gaps from $\mathrm{adv}(f_{\mathrm{adv}})$ ($+24.4$ and $+13.0$, respectively), together with the fact that adversarial training is overly-conservative here as well (as it was in $\mathrm{clean}$), give hope that strategic learning will capitalize on its more focused defense (orange).
While the strength of strategic opponents generally decreases with $k$, there is nonetheless considerable variation in strength across opponents within each class: for example, accuracy against the ‘easiest’ 1-hot opponent is very close to clean ($77.3$, vs. $79.8$), whereas for the ‘hardest’ 1-hot opponent it is midway between adversarial and clean ($59.8$), and worse than some 3-hots. This suggests there is value in knowing not only the strength of the opponent (e.g., the number of feasible targets), but also more fine-grained details regarding its possible incentives. We will examine this idea further in Sec. 6.
@@ 5 Strategic training
Conventional adversarial learning protects against the worst-case perturbations within the feasible set. Under Eq. (3), this can be interpreted as seeking robustness against an opponent whose utility is known to be $u^{\mathrm{adv}}$. Alternatively, Eq. (5) frames adversarial learning as protecting against an unknown opponent by optimizing against the worst-case utility in the uncertainty set $\Lambda$. Operationally, these are equivalent, since $u^{\mathrm{adv}}$ is always the worst-case utility in $\Lambda$—irrespective of the data distribution. The advantage of the utilitarian formulation in Eq. (5) is that it provides a means to reduce the uncertainty set in an informed manner and according to the opponent’s incentive structure. This forms the main idea underlying our approach.
@@
which recovers Eq. (6) for $U=\{u\}$. If we instead have probabilistic beliefs over the possible utilities in a form of a distribution $p_{u}$, then the max can be replaced with an expectation. We refer to the expected error term in Eq. (7) (including the max) as strategic error w.r.t. $U$, and to the general goal of optimizing Eq. (7) as strategic-case robustness (or simply strategic robustness).
@@
Hence, while strategic opponents are worse than no opponent at all, they are generally weaker than adversaries. If the opponent’s true utility lies in $U$, then strategic training should provide robustness—whereas clean training does not, and adversarial training does, but is unnecessarily over-conservative. As we will see, this also results in strategic training performing better on clean data than adversarial training. Strategic learning therefore offers a middle ground between the two extremes of clean and adversarial learning, this giving hope that appropriate strategic modeling can prove to be beneficial for learning.
@@ 5.1 Optimization
Consider first the case of training under a single chosen $u$ (Eq. (6)). Note that since utility is of the form $u(y,y^{\prime})$, for each example $(x,y)$ we need only consider the subset of targets $y^{\prime}$ s.t. $u(y,y^{\prime})=1$. If there is only a single target, then as a proxy we can maximize the probability to predict $y^{\prime}$: denoting ${\hat{p}}(y^{\prime}|x)=\operatorname*{softmax}_{y^{\prime}}(f(x+\delta))$, the strategic response is optimized by $\max_{\delta}{\hat{p}}(y^{\prime}|x)$. Note this mirrors the popular proxy for adversarial attacks, $\min_{\delta}{\hat{p}}(y|x)$. If instead we have multiple possible targets $T\subseteq[K]\setminus y$, then since the opponent gains utility from any one of them, we solve:
${\hat{\delta}}_{u}=\max_{\delta}\max_{y^{\prime}\in T}\,{\hat{p}}(y^{\prime}|x)$ (9)
@@
Next, consider the case of training under an uncertainty set $U$ (Eq. (7)). For general $U$, the simplest approach is to in each batch enumerate over all $u\in U$, find the maximizing $u$, and update $\delta_{u}$ as above accordingly. In principal, finding the argmax should increase runtime by a factor $|U|$. However, since attacks decouple over $y$, we can solve $\max_{u}$ independently for each row $u(y,\cdot)$, which can significantly reduce computational costs: for example, if $U$ includes all 1-hot utilities, then despite having $|U|=(K-1)^{K}$, the total number of calls to Eq. (9) is $O(K)$.
@@ 6 Learning under multi-targeted strategic attacks
We now proceed to demonstrate the potential benefits of strategic training (see Sec. 5), discussing also its potential risks and suggesting ways to hedge them. All experimental details are given in Appx. B.
Figure 4: (Left:) Improvement of strategic training (for known $u$) over adversarial training on a sample of 1-hot opponents, showing gains are larger against ‘easier’ opponents. (Right:) Correlation between strategic accuracy and attack semantics. Opponents who benefit from changing the label’s semantic meaning (e.g., $\mathrm{animal}\mapsto\mathrm{vehicle}$) for more classes enable higher gains.
@@ 6.1 The benefits of accurate strategic modeling
Our first experiment aims to show how strategic training with knowledge of the opponent’s utility can improve performance. Focusing on 1-hot opponents, here we train ResNet18 classifiers on CIFAR-10 using strategic training with a known $u$, and compare their strategic accuracy (w.r.t. $u$) to that of an adversarially trained classifier. This allows us to quantify the boost in accuracy gained by correct incentive modelling. To capture the breadth of possible outcomes, we partition all 1-hot opponents into 10 bins by how adversarial training performs against them (as in Fig. 3) and then sample 15 opponents uniformly from each bin, this serving as a proxy for opponent ‘strength’.
@@ Results.
Fig. 4 (left) shows the gain in strategic accuracy from strategic training across all bins: dots represent single utilities $u$, and boxes show quantiles for each bin. As can be seen, accurate strategic training improves performance on the entire range by at least $+6$ accuracy points (on average) across all bins (see inset plot). As opponents becomes ‘weaker’, average gains increase up to $+10$ points.
@@
Since there is considerable variation across opponents in the gains of strategic training (even within bins), it will be useful to understand what contributes to gains being large. Intuitively, we expect strategic training to improve upon adversarial training when the latter unnecessarily defends against a non-target ($u_{yy^{\prime}}=0$) in a way that makes it susceptible to attacks on true targets ($u_{y{\bar{y}}}=1$). Given this, we conjecture that strategic opponents that benefit from hard-to-attack targets will be easier to defend against using strategic training.
But what makes targets ‘easy’ or ‘hard’ to attack? Fig. 4 (right) suggests that this difficulty is inherently tied to the semantics of the learning task. For example, since CIFAR-10 consists of either animal or vehicle classes, semantic attacks (e.g., $\mathrm{animal}\mapsto\mathrm{animal}$) should be easier for the opponent to accomplish than anti-semantic attacks (e.g., $\mathrm{animal}\mapsto\mathrm{vehicle}$). Here we measure for each $u$ the number of semantic pairs, namely |$|\{y:{\bar{y}}\in S(y)\}|$ where ${\bar{y}}$ is the 1-hot target of $y$. Indeed, results show a strong correlation ($-0.84$) between strategic accuracy and the semantic degree of $u$.
Figure 5: Strategic training against single 1-hot opponents of varying strengths. Results show accuracy for increasing risk mitigation ($\epsilon$) for specified and misspecified settings.
@@ 6.2 The risks of imprecise strategic modeling and how to mitigate them
Our previous results suggest that much can be gained from accurate strategic modeling. But what would happen if at test time we were to face an opponent whose utility is different from the one we trained on? A simple way to hedge such risks is to assume that we will face the chosen $u$ with probability $1-\epsilon$, and any other opponent otherwise. This can be optimized by placing weight $1-\epsilon$ on $u$ and splitting the remainder $\epsilon$ mass uniformly between all other $u^{\prime}\in U$ (e.g., all 1-hots). [^2] In practice this is implemented by sampling at each batch $u$ w.p. $1-\epsilon$ and $u^{\prime}\sim\mathrm{uniform}(U)$ w.p. $\epsilon$. As $\epsilon$ grows, we can expect that if $u$ is well-specified then our gains will decrease, but if it is misspecified, then our losses will be smaller. Note that $\epsilon=1$ amounts to providing average-case robustness w.r.t. $U$.
@@ Results.
Maintaining the previous setup, Fig. 5 demonstrates this idea for strategic 1-hot opponents of varying strengths: randomly chosen (showing mean and standard deviation), and easy and hard representatives. As can be seen, well-specified strategic training improves upon adversarial training for all considered opponents (blue), where gains are indeed higher for small $\epsilon$, but even for the hardest opponent still remain positive even at $\epsilon=0.8$. Conversely, when strategic training is misspecified (orange), performance against the worst-case of all 1-hot opponents is worse than adversarial training. And while the gap does become smaller as $\epsilon$ increases, for all opponents the gap remains even at $\epsilon=1$; this is expected, since evaluation is worst-case, whereas training optimizes for average-case.
@@
Evaluation above is worst-case with respect to all possible opponents; hence, the worst-case $u^{\prime}$ at test time can be arbitrarily different from the $u$ used for training. A more lenient evaluation would be to evaluate against the worst-case $u^{\prime}$ from a smaller set that is ‘closer’ to $u$. Given the relation between attack strength and semantics, Fig. 5 also shows accuracy for misspecified $u$ but evaluated against a semantically relevant set (green). For the hard opponent, the effect disappears; note how adversarial performance remains the same, suggesting that the hardest opponent is a member of the semantic set. However, for the easy opponent, not only is there general improvement, but the gap to adversarial training closes quickly, and at $\epsilon\approx 0.7$ moderately reverses. Note also how here accuracy for the adversary is close to that of the correct setting. Together, these suggest that even the mild knowledge of the adversary’s incentives—e.g., whether it is semantic or anti-semantic—should prove useful.

### v3

@@ 4.1 Adversarial attacks as worst-case strategic responses
In Eq. (3), the form for $\delta^{\mathrm{adv}}$ implies that the adversary benefits from any erroneous prediction; thus, an adversarial opponent is in fact a strategic opponent with utility $u=u^{\mathrm{adv}}\triangleq\ell$ plugged into Eq. (4), and so we can interchangeably think of an adversary as a particular strategic opponent with utility $u^{\mathrm{adv}}$. In general, however, strategic opponents can have other, non-adversarial utilities. Assuming utilities are normalized, let $\Lambda=[0,1]^{K\times K}$ denote the universal set of all utility functions. Then for an adversary, since $u^{\mathrm{adv}}_{y{\hat{y}}}=1$ if $y\neq{\hat{y}}$ and 0 otherwise, our first observation is that Eq. (3) can be rewritten as:
@@ 4.2 The price of adversarial robustness against non-adversarial opponents
Fig. 4 shows the performance of an adversarially trained classifier ($f_{\mathrm{adv}}$) evaluated against different opponents: adversarial ($\mathrm{adv}$), strategic ($\mathrm{1{\text{-}}hot}$ and $\mathrm{3{\text{-}}hot}$), and no opponent ($\mathrm{clean}$). Against $\mathrm{adv}$ (dark red), for which $f_{\mathrm{adv}}$ is consistent in training, accuracy is $45.5$. This increases to $79.8$ in $\mathrm{clean}$ (green over red). This is an improvement—but in comparison, a model trained on clean data ($f_{\mathrm{cln}}$) achieves $93.1$ clean accuracy (green). The price $f_{\mathrm{adv}}$ pays for being conservative is therefore quite steep ($-13.3$; red arrow). Under our utilitarian perspective, this large gap is the result of assuming that all attacks are valuable for the opponent, when in effect none are.
Figure 4: Accuracy of an adversarially trained model against varying opponents: adversarial, clean, and strategic (1-hots and 3-hots), vs. clean accuracy of a clean model. Strategic attacks are harmful, but less than adversarial—a gap which strategic training aims to capitalize on (dashed).
@@
Next, consider how the same adversarially trained model fairs against strategic opponents. Fig. 4 also shows the distribution of strategic accuracies of $f_{\mathrm{adv}}$ against all $k=1$ (light orange) and a large sample of $k=3$ (dark orange) row-wise $k$-hot opponents. Results show that strategic attacks are indeed generally weaker than adversarial attacks, with average values of $69.9\mathrm{1{\text{-}}hot}(f_{\mathrm{adv}})=$ 69.9 and $58.5\mathrm{3{\text{-}}hot}(f_{\mathrm{adv}})=$ 58.5. These large gaps from $\mathrm{adv}(f_{\mathrm{adv}})$ ($+24.4$ and $+13.0$, respectively), together with the fact that adversarial training is overly-conservative here as well (as it was in $\mathrm{clean}$), give hope that strategic learning will capitalize on its more focused defense (dashed).
While the strength of strategic opponents generally increases with $k$, there is nonetheless considerable variation in strength across opponents within each $k$-class: for example, accuracy against the ‘easiest’ 1-hot opponent is very close to clean ($77.3$, vs. $79.8$), whereas for the ‘hardest’ 1-hot opponent it is midway between adversarial and clean ($59.8$), and worse than some 3-hots. This suggests there is value in knowing not only the strength of the opponent (e.g., the number of feasible targets, namely $k$), but also more details regarding its incentive structure. We will examine this idea further in Sec. 6.
@@ 5 Strategic training
Conventional adversarial learning protects against the worst-case perturbations within the feasible set. Under Eq. (3), this can be interpreted as seeking robustness against an opponent whose utility is known to be $u^{\mathrm{adv}}$. Alternatively, Eq. (5) frames adversarial learning as protecting against an unknown opponent by optimizing against the worst-case utility in the uncertainty set $\Lambda$. Operationally, these are equivalent, since $u^{\mathrm{adv}}$ is always the worst-case utility in $\Lambda$—irrespective of the data distribution. The advantage of the utilitarian formulation in Eq. (5) is that it provides a means to reduce the uncertainty set in a principled manner and according to the opponent’s incentive structure. This forms the main idea underlying our approach.
@@
which recovers Eq. (6) for $U=\{u\}$. If we instead have probabilistic beliefs over the possible utilities in a form of a distribution $p_{u}$, then the max can be replaced with an expectation over $p_{u}$. We refer to the expected error term in Eq. (7) (including the max) as strategic error w.r.t. $U$, and to the general goal of optimizing Eq. (7) as strategic-case robustness (or simply strategic robustness).
@@
In this sense, strategic robustness presents a relaxed notion of robustness compared to adversarial. The key modelling component of Eq. (8) is the uncertainty set $U$: if we begin by setting $U=\Lambda$ and gradually reduce it until $U=\{0\}$, we obtain a continuum of learning objectives that interpolate between seeking highly-robust but non-accurate classifiers, and non-robust but highly accurate ones. Hence, by choosing $U$, the learner can determine the desired operating point along this tradeoff front—a choice that should be made deliberately, on the basis of prior knowledge, and by balancing gains and risks.
@@ 5.1 Optimization
Consider first the case of training under a single chosen $u$ (Eq. (6)). Note that since utility is of the form $u(y,y^{\prime})$, for each example $(x,y)$ we need only consider the subset of targets $y^{\prime}$ s.t. $u(y,y^{\prime})=1$. If there is only a single target, then as a proxy we can maximize the probability to predict $y^{\prime}$: denoting ${\hat{p}}(y^{\prime}|x)=\operatorname*{softmax}_{y^{\prime}}(f(x))$, the strategic response is optimized by $\max_{\delta}{\hat{p}}(y^{\prime}|x+\delta)$. Note this mirrors the popular proxy for adversarial attacks, $\min_{\delta}{\hat{p}}(y|x+\delta)$. If instead we have multiple possible targets $T\subseteq[K]\setminus y$, then since the opponent gains utility from any one of them, we solve:
${\hat{\delta}}_{u}=\operatorname*{argmax}_{\delta}\max_{y^{\prime}\in T}\,{\hat{p}}(y^{\prime}|x+\delta)$ (9)
using gradient ascent on $\delta$. For stability, we found it useful to add a small degree of noise: for each example, with probability $\epsilon=0.1$ we replace ${\hat{\delta}}_{u}$ with a targeted attack ${\hat{\delta}}_{y^{\prime}}$ on a random target $y^{\prime}$.
@@
Next, consider the case of training under an uncertainty set $U$ (Eq. (7)). For general $U$, the simplest approach is to in each batch enumerate over all $u\in U$, find the maximizing $u$, and update $\delta_{u}$ as above accordingly. In principle, finding the argmax should increase runtime by a factor $|U|$. However, since attacks decouple over $y$, we can solve $\max_{u}$ independently for each row $u(y,\cdot)$, which can significantly reduce computational costs: for example, if $U$ includes all 1-hot utilities, then despite having $|U|=(K-1)^{K}$, the total number of calls to Eq. (9) is $O(K)$.
@@ 6 Learning under multi-targeted strategic attacks
We now proceed to demonstrate the potential benefits of strategic training (see Sec. 5), discussing also its potential risks and suggesting ways to hedge them. Full experimental details are provided in Appendix B. Extended results on additional datasets and experimental settings are given in Appendix C.
[…]
Figure 5: (Left:) Improvement of strategic training (for known $u$) over adversarial training on a sample of 1-hot opponents, showing gains are larger against ‘easier’ opponents. (Right:) Correlation between strategic accuracy and attack semantics. The more an opponent benefits from predictions that change label semantics (e.g., $\mathrm{animal}\mapsto\mathrm{vehicle}$), the higher the gains of strategic training.
Figure 6: Relative performance of strategic training on random $k$-hot vs. adversarial opponents for increasing $k$. Gains are larger (though variable) for smaller $k$, but remain significant also for large $k$.
@@ 6.1 The benefits of accurate strategic modeling
Our next experiment aims to show how strategic training can use knowledge of the opponent’s utility to improve performance. Focusing on $k$-hot opponents, here we train ResNet18 classifiers on CIFAR-10 using strategic training with a known $u$, and compare their strategic accuracy (w.r.t. $u$) to an adversarially trained classifier. This quantifies the boost in accuracy due to correct incentive modelling.
@@ Results.
For $k=1$, to capture the breadth of possible outcomes, we partition all 1-hot opponents into 10 bins by how adversarial training performs against them (as in Fig. 4) and then sample 15 opponents uniformly from each bin, this serving as a proxy for opponent ‘strength’. Fig. 6 (left) shows the gain in strategic accuracy from strategic training across all bins: dots represent single utilities $u$, and boxes show quantiles for each bin. As can be seen, accurate strategic training improves performance on the entire range by at least $+6$ accuracy points (on average) across all bins (see inset plot). As opponents become ‘weaker’, average gains increase by up to $+10$ points. Fig. 6 shows the relative gains of strategic learning for $k>1$. As expected, gains are large for small $k$, but vary considerably by instance. For large $k$, gains are smaller (and less variable), though still significant ($+3$).
@@
Since there is considerable variation across opponents in the gains of strategic training (even within bins for $k=1$), it will be useful to understand what contributes to gains being large. Intuitively, we expect strategic training to improve upon adversarial training when the latter unnecessarily defends against a non-target ($u_{yy^{\prime}}=0$) in a way that makes it susceptible to attacks on true targets ($u_{y{\bar{y}}}=1$). Given this, we conjecture that strategic opponents that benefit from hard-to-attack targets will be easier to defend against using strategic training.
But what makes targets ‘easy’ or ‘hard’ to attack? Empirically, we observe that a determining factor is the semantic similarity of the target ${\bar{y}}$ and the true label $y$. For CIFAR-10, which comprises animal and vehicle classes, we call a target ${\bar{y}}$ semantic if it is of the same category as $y$ (e.g., $\mathrm{dog}\mapsto\mathrm{cat}$) and anti-semantic if it is not (e.g., $\mathrm{dog}\mapsto\mathrm{truck}$). We then measure for each 1-hot $u$ the number of semantic pairs, namely |$|\{y:{\bar{y}}\in S(y)\}|$ where ${\bar{y}}$ is the target of $y$ in $u$. Fig. 6 (right) shows that accuracy and the number of semantic pairs have a strong negative correlation ($-0.84$). Similar results hold for GTSRB and CIFAR-100 for appropriately defined semantic structures. As we will see, semantics will play an important role in our experimental analysis in the following sections.
Figure 7: Strategic accuracy of strategic training with increasing risk mitigation ($\epsilon$), compared to adversarial training in the same setting. Results are shown for random (mean and std. dev.), easy, and hard 1-hot opponents. For each, performance is evaluated in (i) a well-specified setting where the test-time utility matches the utility used in training, and (ii) a misspecified setting in which test-time utility is set to the worst-case utility (in terms of accuracy) relative to train-time utility. For the latter, results show performance against the worst-case utilities out of all opponents (orange), and w.r.t. a relevant subset of similar opponents (green), either semantic (for hard) or anti-semantic (for easy).
@@ 6.2 The risks of imprecise strategic modeling and how to mitigate them
Our previous results suggest that much can be gained from accurate strategic modeling. But what would happen if at test time we were to face an opponent whose utility is different from the one we trained on? A simple way to hedge such risks is to assume that we will face the chosen $u$ with probability $1-\epsilon$, and any other opponent otherwise. This can be optimized by placing weight $1-\epsilon$ on $u$ and splitting the remainder $\epsilon$ mass uniformly between all other $u^{\prime}\in U$ (e.g., all 1-hots). [^2] In practice this is implemented by sampling at each batch $u$ w.p. $1-\epsilon$ and $u^{\prime}\sim\mathrm{uniform}(U)$ w.p. $\epsilon$. As $\epsilon$ grows, we can expect that if $u$ is well-specified then our gains will decrease, but if it is misspecified, then our losses will be smaller. Note that $\epsilon=1$ amounts to providing average-case robustness w.r.t. $U$. We present results for this approach under both correct and false assumptions on the opponent’s incentives.
@@ Results.
Fig. 7 shows results for random 1-hot opponents (left) as well as focused results against the easiest (center) and hardest (right) opponent, taken from Sec. 4.2. In the well-specified setting where performance is evaluated on the same $u$ used in training (blue), strategic training improves upon adversarial training in all cases. As expected, gains decrease as $\epsilon$ increases, but remain positive even against the hardest opponent and at $\epsilon=0.8$. For the misspecified setting, we show the lowest accuracy obtained against any 1-hot utility $u^{\prime}\neq u$ (orange). Results show that increasing $\epsilon$ provides increased protection against misspecification, though even for $\epsilon=1$ it offers less than adversarial training. For $k>1$, results show similar trends but with lower overall accuracy for larger $k$ (see Appendix C.2).
@@
The risks of misspecification were measured above with respect to the worst-case out of all possible opponents. This considers a demanding setting in which $u^{\prime}$ at test time can be arbitrarily different from the $u$ used for training. A more lenient evaluation would be to consider $u^{\prime}$ that is still worst-case, but out of a smaller set that is ‘closer’ to $u$. Given the relation between attack strength and semantics, Fig. 7 also shows accuracy under misspecification against the worst-case but semantically relevant opponent (green): anti-semantic for the easy opponent, and semantic for the hard opponent. For the latter, note how the effect of misspecification disappears; since adversarial performance remains the same, this suggests that the hardest opponent is indeed a member of the semantic set. However, for the easy opponent, not only is there general improvement, but the gap to adversarial training closes quickly, and at $\epsilon\approx 0.7$ moderately reverses. Note also how here accuracy for the adversary is close to that of the correct setting. Together, these suggest that even the mild knowledge of the adversary’s incentives—e.g., whether it is semantic or anti-semantic—can be useful.